please help me solve this question-I have three children. The youngest is the same age as the first digit in my age another is the same age as the second digit in my age, and the third is the same age as the sum of the two digits in my age. The total of our ages is 33. How old am I?
Answer by ankor@dixie-net.com(22740)

I have three children.
The youngest is the same age as the first digit in my age
another is the same age as the second digit in my age,
and the third is the same age as the sum e two digits in my age.
The total of our ages is 33. How old am I?
:
let a = the 1st digit of your age
let b = the 2nd digit
then
10a+b = your age
:
"third is the same age as the sum 0f the two digits in my age"
a+b = third child's age
:
"The total of our ages is 33"
10a+b + a + b + a+b = 33
12a + 3b = 33
simplify, divide by 3
4a + b = 11
b = -4a + 11
From this, we can see a has to be 1 or 2, b can't be a neg number
Also a has to be 2 since you are an adult and the total is 33
So we have:
a = 2 yrs 1st child
b = 3 yrs 2nd child
a+b= 5 yrs 3rd child
2(10)+3 = 23 yrs is your age
--------------
total is 33